OpenStudy (cherie_magee):

1/2x + 1/3y = 0 1/4x - 1/2y = 8 What is the solution of the system

OpenStudy (anonymous):

Have u tried it ??

OpenStudy (cherie_magee):

yes i have but i keep getting confused @Anu2401

OpenStudy (anonymous):

You can use substitution method . Just find the value of x from any one of the equation & put this value of x(Which will be in terms of y) into the other equation

OpenStudy (cherie_magee):

that just sounded likw another lingo lol @Anu2401

OpenStudy (anonymous):

Ok i"ll give u a hint . Have a look @ attachment ... Try to proceed from here

OpenStudy (anonymous):

@cherie_magee: Can u proceed from here to solve the question ??

OpenStudy (anonymous):

See it When u r done ........ Njoy

OpenStudy (texaschic101):

you can make this easier by multiplying the first equation by 6 and the second equation by 4 and this will get rid of the fractions

OpenStudy (mathstudent55):

@Anu2401 is solving this system of equations: \( \dfrac{1}{2x} + \dfrac{1}{3y} = 0 \) \( \dfrac{1}{4x} - \dfrac{1}{2y} = 8 \) Is that correct? Is the system what Anu thinks it is or is it (which @texaschic101 is solving)? \( \dfrac{1}{2}x + \dfrac{1}{3}y = 0 \) \( \dfrac{1}{4}x - \dfrac{1}{2}y = 8 \)

OpenStudy (anonymous):

@mathstudent55 : Thanks for pointing it out bro :) Very nice analysis :)

OpenStudy (texaschic101):

so which set of equations is it ?

OpenStudy (mathstudent55):

My guess is it's the second one, but it can be either.

OpenStudy (texaschic101):

well....I suppose when or if she responds, we will know......lol